3. There is a commercial on TV right now that says many people spend more money on their morning Starbucks than they invest in their retirement savings. Lets say you stop at Starbucks every morning on the way to school and buy a $6 cup of fancy coffee. Assuming, for simplicity, that there are 30 days in a month, thats $180 per month you could be investing in your retirement. If you deposit $180 per month into a savings account paying 0.25% per month compound interest, how much will you have saved up after 40 years of working as an engineer?
Answer #1

ANSWER:

n = 40 years or 40 * 12 = 480 months

i = 0.25%

monthly savings = $180

savings after 40 years = monthly savings(p/a,i,n)

savings after 40 years = 180(p/a,0.25%,480)

savings after 40 years = 180 * 279.3

savings after 40 years = 50,274

so after 40 years the savings would be $50,274